Exercise 9-25

Marigold Company began operations late in 2016 and adopted the conventional retail inventory method. Because there was no beginning inventory for 2016 and no markdowns during 2016, the ending inventory for 2016 was $13,737 under both the conventional retail method and the LIFO retail method. At the end of 2017, management wants to compare the results of applying the conventional and LIFO retail methods. There was no change in the price level during 2017. The following data are available for computations.

Cost

Retail

Inventory, January 1, 2017 $13,737 $19,500
Sales revenue 87,000
Net markups 8,600
Net markdowns 1,500
Purchases 63,300 88,000
Freight-in 14,682
Estimated theft 1,800


Compute the cost of the 2017 ending inventory under both:

(a) The conventional retail method. (Round ratios for computational purposes to 0 decimal places, e.g. 78% and final answer to 0 decimal places, e.g. 28,987.)

Ending inventory using the conventional retail method $


(b) The LIFO retail method. (Round ratios for computational purposes to 0 decimal places, e.g. 78% and final answers to 0 decimal places, e.g. 28,987.)

Ending inventory at cost $

Ending inventory at retail $
Answer #1

ANSWER

1)conventional retail method:

Cost Retail
Beginning 13,737 19,500
Purchase 63,300 88,000
Freight in 14,682
Mark up 8,600
Inventory after markup 91,719 116,000
less:markdown -1,500
Inventory after markdown 114,600
less:sales -87,000
Estimated theft -1,800
Estimated inventory at retail 25,800

cost to retail ratio = 91,719/116,000= .79 or 79%

Ending inventory at Cost : cost to retail ratio * estimated inventory at retail'              

                    = .73*25,800

                   = $ 20,382

2)

cost Retail cost to retail ratio
Beginning 13,737 19,500 13,737/19,500= .70 or 70%
Purchase 63,300 88,000
Freight in 14,682
Mark up 8,600
less:mark down -1,500
Inventory without beginning inventory 77,982 95,100 77,982/95,100=.82 or 82%
Inventory with beginning inventory 91,719 114,600
less:sales -87,000
Estimated theft -1,800
Inventory at retail 25,800[114,600-87,000-1,800]

under LIFO units acquired last are sold first so ending inventory are left from beginning inventory.

cost Retail
Ending inventory at cost left from 2016 (beginning) 19,500*.70= 13,650 19,500
Ending inventory at cost left from 2017 6300*.82=5,166 25,800-19,500= 6,300
Total 18816 25,800

_____________________________________________

If you have any query or any Explanation please ask me in the comment box, i am here to helps you.please give me positive rating.

*****************THANK YOU**************